OpenStudy (anonymous):

Cos(theta)=(15/17), and sin>0; How do I solve for 2theta?

OpenStudy (anonymous):

Use the trig identities: a) cos 2a = 2*cos^2 a -1. Replace cos a by 15/17 cos 2a = 2*(225/289) - 1 = 1.58 - 1 = 0.58 b) sin^2 a + cos^2 a = 1 sin^2 a = 1 - cos^2 a= 1 - (225/289) = 1 - 0.78 = 0.22 sin a = V(0.22) = 0.47. c) sin 2a = 2*sin a*cos a = 2(0.47)(15/17) = ......